OpenStudy (anonymous):

need help finding the slope of x+5y=10

jimthompson5910 (jim_thompson5910):

x+5y=10 5y=-x+10 y=-1/5x+10/5 y=-1/5x+2 So the slope is -1/5

OpenStudy (anonymous):

convert to y=mx+c m is the slope